OpenStudy (anonymous):

HELP ME PLEASEEEE!!! - A bakery charges $0.30 for a cookie that is 3 inches in diameter. If the price is proportional to the area, how much do they charge for an extra large cookie that is 6 inches in diameter? Assume that each cookie is shaped of a circle.

OpenStudy (anonymous):

can you help me, and i will help you back as well?

OpenStudy (anonymous):

what do you needl help on?

OpenStudy (anonymous):

A = πr^2 A = π(9/4) A = 2.25π 0.3 : 2.25π = x : 9π x = 1.20π or x ≈ 3.77

OpenStudy (anonymous):

What are the x-intercept(s) of the graph of the equation y − x^2 = −4x − 5? (1, 0) and (−5, 0) (1, 5) (−1, 0) and (−5, 0) (−1, 0) and (5, 0)

OpenStudy (anonymous):

To find the x intercepts, set y = 0, and solve. BTW, note that answer B cannot be correct because it doesn't have y = 0! So with y=0 you have ---> x2 - 4x - 5 = 0 This can be expressed as (x+1)*(x-5) = 0 This can be solved separately as x+1 = 0 and x-5 = 0 to find the two answers. the answer is D